# Determining the optimal structural resolution of proteins through an information-theoretic analysis of their conformational ensemble

### v5

@@ Abstract
American Chemical Society,
@@ 1 Introduction
It has to be acknowledged, however, that for the sake of the analysis an extremely rich description of the system is not always necessary nor desirable: indeed, an excess of detail can obfuscate important large-scale features that are key to grasp the essential characteristics of the system. As an extreme example, the precise and exact knowledge of each atom of a protein is not useful in determining the relative population of two conformational basins; instead, fewer (even a single one) and much coarser variables might suffice 13, 14, 15. This illustrates the important difference between data and information: one thing is to have many distinct configurations, another thing is to group them together and see states emerge.
@@
In computer simulations, this issue appears vigorously. Long, time-intensive molecular dynamics (MD) trajectories are necessary to achieve an adequately extensive coverage of the configurational space of a protein, and this need becomes especially painstaking as the size of the molecule increases. At the same time, though, the enormous number of distinct frames generated can be redundant with respect to certain information one might aim for. In a previous work 16, some of us have addressed this issue by grouping MD frames through a hierarchical clustering approach: varying the distance threshold below which two frames are grouped together, one can look for the level of configurational space coarsening that attains a compromise between data parsimony (smallest number of clusters) and informativeness (each cluster is representative of a distinct state). It was shown that an optimal trade-off point indeed exists, at which a large frame dataset can be compressed not only without losing important detail, but even revealing significant features.
@@
The aforementioned work relied on a structure-based clustering of the protein conformations, in which either all heavy atoms or all Cα atoms were retained. Yet, the idea that the highest level of detail is not necessarily the most useful applies not only to the sampled state space, but rather also to the structure itself. In fact, while all atoms in the protein contribute to its emergent behaviour, not all of them contribute equally, because of the numerous internal constraints and restraints (chemical bonds, non-bonded interactions, free energy basins…) that limit the conformational variability of the degrees of freedom. Furthermore, the very existence and the successful application of coarse-grained models 17, 18, 19, 20 demonstrate that a simplified structural representation of proteins can suffice to reproduce, qualitatively and sometimes even quantitatively, many emergent properties that occur at a typical scale larger than that of the groups of atoms lumped into effective interaction sites. However, while it is relatively easy to embrace the idea that a low-resolution representation of a protein can be useful and informative, finding such a representation is a much more complex endeavour. The question thus arises, whether it is possible to quantify in an unambiguous manner the informativeness of a coarse-grained representation of a protein and, by maximising it, to identify an optimal resolution level, that is, the optimal number of atoms to be preserved in the low-resolution representation.
@@
In practice, we here address this problem as follows. First, as coarse-grained representations we consider decimation mappings 21, 22 of a protein, that is, selections of a subset of its heavy atoms, and we investigate the properties of the molecule in terms of the number of such atoms. Second, given an MD simulation of the protein, we cluster the sampled conformations based on the selected subset, and repeat the clustering for many different atoms at the same resolution level. Lastly, we quantify the informativeness of each clustering and search for the resolution level that, on average, maximises the retained information while minimising the detail.
Key to this last step is evidently a quantitative notion of informativeness. There is no unique way in which one can measure the amount of information entailed by a dataset, and the most adequate approach is generally problem-specific. Here, it is our objective to establish how informative a given clustering is about the generative process that underlies the sampled MD frames, implying, by this, that the statistical weights of the clusters reverberate the depths of the associated free energy basins. To do so, we employ critical variable selection 23, 24, an information theory approach that quantifies, with distinct metrics, the level of detail of the representation of a given ensemble (its resolution) and the amount of useful information it contains about the stochastic process that generated the data (the relevance).
@@ 2 The PROPRE method: theoretical background and algorithmic aspects
The PROPRE approach consists in a protocol for the analysis of an atomistic trajectory of a biomolecule, aimed at identifying the optimal number $N_{opt}$ of heavy atoms to retain in a synthetic yet informative low-resolution description of the system and its conformational space. In other words, the method allows the determination of the optimal level of resolution to capture the essential features of the biomolecule’s large-scale dynamics, retaining the smallest amount of detail while preserving the largest amount of information.
@@ 2.1 Critical variable selection, or resolution and relevance framework
The PROPRE protocol builds on the critical variable selection (CVS), or resolution–relevance framework, introduced by Marsili and co-workers 25, 26, 24 and applied across diverse domains, from protein sequences 23 to neural activity 27 and protein conformational ensembles 16. This approach provides an information-theoretic criterion to identify informative coarse representations of complex datasets. Detailed theoretical discussions can be found elsewhere 25, 26, 24, 16, 28; here we summarize the key definitions and their role within PROPRE workflow.
@@
where $k_{s}$ is the number of elements in cluster $s$. Resolution quantifies the level of detail, ranging from $H_{\mathrm{res}}=0$ when all elements belong to a single group to $H_{\mathrm{res}}=\log M$ when every element is distinct. The complementary notion of relevance captures how informative the clustering is, by measuring the heterogeneity in cluster sizes 25, 24:
@@
Relevance vanishes at both extremes of resolution, since both the trivial partition (all elements in one cluster) and the maximally detailed one (all singletons) evidence no statistical structure. Between these extremes, $H_{\mathrm{rel}}[k]$ attains a maximum, which identifies the clustering that retains the largest amount of non-redundant information about the data. The relation between resolution and relevance takes the form of a bell-shaped curve, whose interpretation is central to the framework. The maximum of the curve corresponds to the most informative compression of the dataset, where heterogeneity in cluster occupancies is most pronounced ($N_{\mathrm{OPT}}^{\mathrm{MR}}$). Equally important is the point where the slope of the curve equals $\mu=-1$ ($N_{\mathrm{OPT}}^{\mathrm{IT}}$), which signals an optimal trade-off: here a unit loss in resolution is exactly balanced by a unit gain in informativeness 25, 24. From an information-theoretic perspective, these two points highlight the most meaningful coarse-grained descriptions of the dataset, either by maximizing relevance or by ensuring an optimal balance between detail loss and information gain.
@@ 2.2 Calculation of the resolution-relevance curves
The algorithm scans a range of numbers of retained (heavy) atoms, denoted $N_{\mathrm{CG}}$, and computes, for each value, the corresponding resolution and relevance over a sample of randomly generated mappings. In this way, we can construct a resolution-relevance curve in which the optimal resolution value, associated with the slope point 1-1, can be identified. A schematic illustration of the whole workflow is provided in Fig. 2.
@@
As anticipated, the starting point of the PROPRE algorithm is the molecular dynamics trajectory of a biomolecule, which contains all atoms of the system of interest but not those of the solvent. Denoting with $n_{\mathrm{ha}}$ the number of heavy atoms in the molecule, the atomistic trajectory $\mathbf{X}_{\mathrm{ha}}(t)$ consists of a collection of $T$ vectors of dimension $3n_{\mathrm{ha}}$, each labelled with the time index $t$ associated with the given time frame. In the applications presented here, we use only the heavy-atom coordinates; other non–system-specific features could be incorporated without affecting its generality or unsupervised character.
The UPGMA 29 hierarchical clustering algorithm is then applied to cluster the configurations of $\mathbf{X}_{\mathrm{ha}}(t)$. The metric employed to measure the distance between two configurations $\mathbf{X}^{a}$ and $\mathbf{X}^{b}$ is the root squared distance (RSD), defined as:
@@
The RSD is proportional to the more common root mean squared distance, $\mathrm{RSD}=\sqrt{n_{\mathrm{ha}}}\cdot\mathrm{RMSD}$, which, despite known limitations 30, 31, still represents the most reliable, interpretable, and commonly used measure to quantify the dissimilarity between two protein structures 32.
@@
The choice of employing the RSD is motivated by advantages it entails in the present context 21, 28. Specifically, if we employed the RMSD as a measure of distance between pairs of configurations, we would have to set, for each value of $N_{\mathrm{CG}}$, a different clustering threshold, since the RMSD is normalized by the number of sites. By contrast, the usage of the RSD preserves the absolute scale of distances, allowing us to choose a single threshold that can be employed irrespective of the number of sites retained at each iteration step. The aforementioned RSD distance threshold $\tilde{C}$ is the minimal value of dissimilarity, that is, the smallest distance below which two configurations can be considered to be structurally identical. To set this value, we perform a hierarchical clustering of the all-atom trajectory, and define $\tilde{C}$ as the largest value below which all frames are distinguishable, i.e. each frame is representative of a singleton cluster. The threshold $\tilde{C}$ is then employed in the next step of the workflow; here we consider a number $N_{\mathrm{CG}}^{(i)}=n_{\mathrm{ha}}-i$ of retained atoms for $i\in[2,n_{\mathrm{ha}}]$. A number $\mathcal{N}_{M}$ of different mappings is generated, that is, a random selection of $N_{\mathrm{CG}}^{(i)}$ heavy atoms out of the $n_{\mathrm{ha}}$ available ones. For each one of the $\mathcal{N}_{M}$ mappings, the trajectory $\mathbf{X}_{\mathrm{ha}}(t)$ is processed to obtain the CG configuration $\mathbf{X}^{(m,i)}_{\mathrm{CG}}(t)$ of the sole atoms that are included in the mapping, the index $m$ labelling the specific mapping and $i$ indicating the retained atom number $N_{\mathrm{CG}}^{(i)}$. The frames of this trajectory are clustered with the UPGMA algorithm, and the dendrogram is cut at the threshold value $\tilde{C}$: in this way, two CG configurations are considered indistinguishable if their distance falls below the same threshold used to resolve structural differences in the all-atom space. This ensures a consistent resolution criterion across both representations: CG configurations are grouped together only if the retained structural differences would have been negligible at full resolution, and treated as distinct only if their full-atom counterparts were also distinguishable.
@@
Figure 3: Plot of resolution ($H_{\mathrm{res}}$) vs. relevance ($H_{\mathrm{rel}}$) obtained for a trajectory of the enzyme adenylate kinase. Points are coloured according to the number of atoms retained in the corresponding mapping $N_{\mathrm{CG}}$. The black dashed lines represent the upper and lower bound to the theoretical maximum, derived in 25, 33. The grey dotted line shows the typical behaviour of a structure-less sample obtained by averaging over multiple random partitions of M balls in an increasing number of boxes.
@@
The clusters so obtained are then used to calculate one $(H_{\mathrm{res}},H_{\mathrm{rel}})$ point; see Eqs. 1 and 2. This procedure is repeated for each choice of the number of retained atoms $N_{\mathrm{CG}}^{(i)}$ and for each random mapping $m$ generated retaining $N_{\mathrm{CG}}^{(i)}$ atoms. The resulting points draw a curve in the resolution-relevance plane as the one shown in Fig. 3 for the protein adenylate kinase (see below).
The resulting curve can then be analysed to extract characteristic reference values ($N_{\mathrm{OPT}}^{\mathrm{IT}}$ and $N_{\mathrm{OPT}}^{\mathrm{MR}}$). The practical procedure used to identify them, including smoothing and numerical differentiation, is described in the Supplementary Information (SI).
@@ 3.1 Protein selection
A dataset of 11 proteins distinct in size, structure and function was built starting from a larger dataset of 107 proteins, after clusterization based on their dynamics. The initial, large dataset was based on the one employed in reference 34, from which proteins with too high sequence/structure similarity were removed. The proteins are categorized monomeric by the protein quaternary structure file server (PQS), have a resolution higher than 1.8Å, no ligands larger than six atoms, and no presence of metals other than Mg2+, Ca2+, K+, Na+, Zn+. Further requirements included a structure deposition date after 1987 and absence of gaps larger than one amino acid. The structures were downloaded from the Protein Data Bank, and the coordinate files were cleaned-up from heteroatoms, from copies of the protein in the crystallographic cell, and from residue-configurations with low occupancy. The position of missing atoms was rebuilt and the protein conformations were optimized using the software FoldX 35. The procedure of dynamics-based clustering applied on this dataset reflects the one developed by us in 36, and already applied in 16. Specifically, for each protein of the larger dataset, the first 10 normal modes of fluctuation were analysed using an elastic network model, and superimposed by means of the ALADYN protocol 37, which performs a hybrid structural/dynamical alignment. The similarity between the essential spaces spanned by the first 10 normal modes was quantified by means of the root mean square inner product (RMSIP) 38. The distance between the essential dynamics of two aligned proteins was defined as $d_{ij}=1-\textrm{RMSIP}_{ij}$; this distance was employed to perform a hierarchical clustering. The 11 proteins used in this work are the centroids of the resulting clusters, and their PDB codes are: 1DSL, 1NOA, 1SNO, 1UNE, 1XWL, 1IGD, 1HYP, 1KNT, 1QKE, 2EXO, 1KOE.

### v6

@@ Abstract
Information theory, Coarse-graining, Multiscale modelling, Protein conformational ensembles, Molecular dynamics simulations, Optimal structural resolution
@@ 1 Introduction
It has to be acknowledged, however, that for the sake of the analysis an extremely rich description of the system is not always necessary nor desirable: indeed, an excess of detail can obfuscate important large-scale features that are key to grasp the essential characteristics of the system. As an extreme example, the precise and exact knowledge of each atom of a protein is not useful in determining the relative population of two conformational basins; instead, fewer (even a single one) and much coarser variables might suffice 13, 14, 15. This illustrates the important difference between data and information: one thing is to have many distinct configurations, another thing is to group them together and see states emerge 16, 17.
@@
In computer simulations, this issue appears vigorously. Long, time-intensive molecular dynamics (MD) trajectories are necessary to achieve an adequately extensive coverage of the configurational space of a protein, and this need becomes especially painstaking as the size of the molecule increases. At the same time, though, the enormous number of distinct frames generated can be redundant with respect to certain information one might aim for. In a previous work 18, some of us have addressed this issue by grouping MD frames through a hierarchical clustering approach: varying the distance threshold below which two frames are grouped together, one can look for the level of configurational space coarsening that attains a compromise between data parsimony (smallest number of clusters) and informativeness (each cluster is representative of a distinct state). It was shown that an optimal trade-off point indeed exists, at which a large frame dataset can be compressed not only without losing important detail, but even revealing significant features.
@@
The aforementioned work relied on a structure-based clustering of the protein conformations, in which either all heavy atoms or all Cα atoms were retained. Yet, the idea that the highest level of detail is not necessarily the most useful applies not only to the sampled state space, but rather also to the structure itself. In fact, while all atoms in the protein contribute to its emergent behaviour, not all of them contribute equally, because of the numerous internal constraints and restraints (chemical bonds, non-bonded interactions, free energy basins…) that limit the conformational variability of the degrees of freedom. Furthermore, the very existence and the successful application of coarse-grained models 19, 20, 21, 22 demonstrate that a simplified structural representation of proteins can suffice to reproduce, qualitatively and sometimes even quantitatively, many emergent properties that occur at a typical scale larger than that of the groups of atoms lumped into effective interaction sites. However, while it is relatively easy to embrace the idea that a low-resolution representation of a protein can be useful and informative, finding such a representation is a much more complex endeavour. The question thus arises, whether it is possible to quantify in an unambiguous manner the informativeness of a coarse-grained representation of a protein and, by maximising it, to identify an optimal resolution level, that is, the optimal number of atoms to be preserved in the low-resolution representation.
@@
In practice, we here address this problem as follows. First, as coarse-grained representations we consider decimation mappings 23, 24 of a protein, that is, selections of a subset of its heavy atoms, and we investigate the properties of the molecule in terms of the number of such atoms. Second, given an MD simulation of the protein, we cluster the sampled conformations based on the selected subset, and repeat the clustering for many different atoms at the same resolution level. Lastly, we quantify the informativeness of each clustering and search for the resolution level that, on average, maximises the retained information while minimising the detail.
Key to this last step is evidently a quantitative notion of informativeness. There is no unique way in which one can measure the amount of information entailed by a dataset, and the most adequate approach is generally problem-specific. Here, it is our objective to establish how informative a given clustering is about the generative process that underlies the sampled MD frames, implying, by this, that the statistical weights of the clusters reverberate the depths of the associated free energy basins. To do so, we employ critical variable selection 25, 26, an information theory approach that quantifies, with distinct metrics, the level of detail of the representation of a given ensemble (its resolution) and the amount of useful information it contains about the stochastic process that generated the data (the relevance).
@@ 2 The PROPRE method: theoretical background and algorithmic aspects
The PROPRE approach consists in a protocol for the analysis of an atomistic trajectory of a biomolecule, aimed at identifying the optimal number $N_{\mathrm{OPT}}$ of heavy atoms to retain in a synthetic yet informative low-resolution description of the system and its conformational space. In other words, the method allows the determination of the optimal level of resolution to capture the essential features of the biomolecule’s large-scale dynamics, retaining the smallest amount of detail while preserving the largest amount of information.
@@ 2.1 Critical variable selection, or resolution and relevance framework
The PROPRE protocol builds on the critical variable selection (CVS), or resolution-relevance framework, introduced by Marsili and co-workers 27, 28, 26 and applied across diverse domains, from protein sequences 25 to neural activity 29 and protein conformational ensembles 18. This approach provides an information-theoretic criterion to identify informative coarse representations of complex datasets. Detailed theoretical discussions can be found elsewhere 27, 28, 26, 18, 30; here we summarize the key definitions and their role within PROPRE workflow.
@@
where $k_{s}$ is the number of elements in cluster $s$. Resolution quantifies the level of detail, ranging from $H_{\mathrm{res}}=0$ when all elements belong to a single group to $H_{\mathrm{res}}=\log M$ when every element is distinct. The complementary notion of relevance captures how informative the clustering is, by measuring the heterogeneity in cluster sizes 27, 26:
@@
Relevance vanishes at both extremes of resolution, since both the trivial partition (all elements in one cluster) and the maximally detailed one (all singletons) evidence no statistical structure. Between these extremes, $H_{\mathrm{rel}}[k]$ attains a maximum, which identifies the clustering that retains the largest amount of non-redundant information about the data. The relation between resolution and relevance takes the form of a bell-shaped curve, whose interpretation is central to the framework. The maximum of the curve corresponds to the most informative compression of the dataset, where heterogeneity in cluster occupancies is most pronounced ($N_{\mathrm{OPT}}^{\mathrm{MR}}$). Equally important is the point where the slope of the curve equals $\mu=-1$ ($N_{\mathrm{OPT}}^{\mathrm{IT}}$), which signals an optimal trade-off: here a unit loss in resolution is exactly balanced by a unit gain in informativeness 27, 26. From an information-theoretic perspective, these two points highlight the most meaningful coarse-grained descriptions of the dataset, either by maximizing relevance or by ensuring an optimal balance between detail loss and information gain.
@@ 2.2 Calculation of the resolution-relevance curves
The algorithm scans a range of numbers of retained (heavy) atoms, denoted $N_{\mathrm{CG}}$, and computes, for each value, the corresponding resolution and relevance over a sample of randomly generated mappings. A schematic illustration of the whole workflow is provided in Fig. 2.
@@
As anticipated, the starting point of the PROPRE algorithm is the molecular dynamics trajectory of a biomolecule, which contains all atoms of the system of interest but not those of the solvent. Denoting with $n_{\mathrm{ha}}$ the number of heavy atoms in the molecule, the atomistic trajectory $\mathbf{X}_{\mathrm{ha}}(t)$ consists of a collection of $T$ vectors of dimension $3n_{\mathrm{ha}}$, each labelled with the time index $t$ associated with the given time frame. In the applications presented here, we use only the heavy-atom coordinates; other non-system-specific features could be incorporated without affecting its generality or unsupervised character.
The UPGMA 31 hierarchical clustering algorithm is then applied to cluster the configurations of $\mathbf{X}_{\mathrm{ha}}(t)$. The metric employed to measure the distance between two configurations $\mathbf{X}^{a}$ and $\mathbf{X}^{b}$ is the root squared distance (RSD), defined as:
@@
The RSD is proportional to the more common root mean squared distance, $\mathrm{RSD}=\sqrt{n_{\mathrm{ha}}}\cdot\mathrm{RMSD}$, which, despite known limitations 32, 33, still represents the most reliable, interpretable, and commonly used measure to quantify the dissimilarity between two protein structures 34.
@@
The choice of employing the RSD is motivated by advantages it entails in the present context 23, 30. Specifically, if we employed the RMSD as a measure of distance between pairs of configurations, we would have to set, for each value of $N_{\mathrm{CG}}$, a different clustering threshold, since the RMSD is normalized by the number of sites. By contrast, the usage of the RSD preserves the absolute scale of distances, allowing us to choose a single threshold that can be employed irrespective of the number of sites retained at each iteration step. The aforementioned RSD distance threshold $\tilde{C}$ is the minimal value of dissimilarity, that is, the smallest distance below which two configurations can be considered to be structurally identical. To set this value, we perform a hierarchical clustering of the all-atom trajectory, and define $\tilde{C}$ as the largest value below which all frames are distinguishable, i.e. each frame is representative of a singleton cluster. The threshold $\tilde{C}$ is then employed in the next step of the workflow; here we consider a number $N_{\mathrm{CG}}^{(i)}=n_{\mathrm{ha}}-i$ of retained atoms for $i\in[2,n_{\mathrm{ha}}]$. A number $\mathcal{N}_{M}$ of different mappings is generated, that is, a random selection of $N_{\mathrm{CG}}^{(i)}$ heavy atoms out of the $n_{\mathrm{ha}}$ available ones. For each one of the $\mathcal{N}_{M}$ mappings, the trajectory $\mathbf{X}_{\mathrm{ha}}(t)$ is processed to obtain the CG configuration $\mathbf{X}^{(m,i)}_{\mathrm{CG}}(t)$ of the sole atoms that are included in the mapping, the index $m$ labelling the specific mapping and $i$ indicating the retained atom number $N_{\mathrm{CG}}^{(i)}$. The frames of this trajectory are clustered with the UPGMA algorithm, and the dendrogram is cut at the threshold value $\tilde{C}$: in this way, two CG configurations are considered indistinguishable if their distance falls below the same threshold used to resolve structural differences in the all-atom space. This ensures a consistent resolution criterion across both representations: CG configurations are grouped together only if the retained structural differences would have been negligible at full resolution, and treated as distinct only if their full-atom counterparts were also distinguishable.
@@
Figure 3: Plot of resolution ($H_{\mathrm{res}}$) vs. relevance ($H_{\mathrm{rel}}$) obtained for a trajectory of the enzyme adenylate kinase. Points are coloured according to the number of atoms retained in the corresponding mapping $N_{\mathrm{CG}}$. The black dashed lines represent the upper and lower bound to the theoretical maximum, derived in 27, 35. The grey dotted line shows the typical behaviour of a structure-less sample obtained by averaging over multiple random partitions of M balls in an increasing number of boxes.
@@
The clusters so obtained are then used to calculate one $(H_{\mathrm{res}},H_{\mathrm{rel}})$ point; see Eqs. 1 and 2. This procedure is repeated for each choice of the number of retained atoms $N_{\mathrm{CG}}^{(i)}$ and for each random mapping $m$ generated retaining $N_{\mathrm{CG}}^{(i)}$ atoms. The resulting points draw a curve in the resolution-relevance plane as the one shown in Fig. 3 for the protein adenylate kinase (see below). These data can then be analysed to identify “special” resolution values, namely the point where the curve has slope 1-1 ($N_{\mathrm{OPT}}^{\mathrm{IT}}$) and the point corresponding to the maximum of the relevance ($N_{\mathrm{OPT}}^{\mathrm{MR}}$). In the following, we will present the results obtained employing both criteria unless otherwise stated; the algorithmic procedure used to identify such values, including smoothing and numerical differentiation of the resolution-relevance curve, is described in the Supplementary Information (SI).
@@ 3.1 Protein selection
A dataset of 11 proteins distinct in size, structure and function was built starting from a larger dataset of 107 proteins, after clusterization based on their dynamics. The initial, large dataset was based on the one employed in reference 36, from which proteins with too high sequence/structure similarity were removed. The proteins are categorized monomeric by the protein quaternary structure file server (PQS), have a resolution higher than 1.8Å, no ligands larger than six atoms, and no presence of metals other than Mg2+, Ca2+, K+, Na+, Zn+. Further requirements included a structure deposition date after 1987 and absence of gaps larger than one amino acid. The structures were downloaded from the Protein Data Bank, and the coordinate files were cleaned-up from heteroatoms, from copies of the protein in the crystallographic cell, and from residue-configurations with low occupancy. The position of missing atoms was rebuilt and the protein conformations were optimized using the software FoldX 37. The procedure of dynamics-based clustering applied on this dataset reflects the one developed by us in 38, and already applied in 18. Specifically, for each protein of the larger dataset, the first 10 normal modes of fluctuation were analysed using an elastic network model, and superimposed by means of the ALADYN protocol 39, which performs a hybrid structural/dynamical alignment. The similarity between the essential spaces spanned by the first 10 normal modes was quantified by means of the root mean square inner product (RMSIP) 40. The distance between the essential dynamics of two aligned proteins was defined as $d_{ij}=1-\textrm{RMSIP}_{ij}$; this distance was employed to perform a hierarchical clustering. The 11 proteins used in this work are the centroids of the resulting clusters, and their PDB codes are: 1DSL, 1NOA, 1SNO, 1UNE, 1XWL, 1IGD, 1HYP, 1KNT, 1QKE, 2EXO, 1KOE.